OpenStudy (anonymous):

someone HELP PLEASE!!! The product of a number and the number increased by 9 is 15. What is the larger of the 2 solutions?

OpenStudy (mertsj):

Why did you post this again?

OpenStudy (anonymous):

because, I can and you didn't help me I got it wrong.

OpenStudy (mertsj):

I did help you and the answer was correct. You didn't enter it correctly.

OpenStudy (anonymous):

how did i not enter it correctly ? how are you supposed to enter it

OpenStudy (mertsj):

Your directions will tell you. Post them.

OpenStudy (anonymous):

Don't include spaces. Such as x + 3 may be marked wrong but x+3 is correct. Don't include symbols. Such as $25 is wrong but 25 is correct. 2x is correct but 2*x or 2(x) are marked wrong. If the question asks for an expression, then = is never used. If the question asks for an equation, then = must be used. Don't round unless stated to do so in the problem. If the answer is a fraction, write the fraction instead of decimal unless the problem states otherwise. Always write in descending order, unless the problem specifies otherwise.

OpenStudy (mertsj):

Don't round unless stated to do so in the problem.

OpenStudy (mertsj):

You rounded.

OpenStudy (anonymous):

I didn't round!!!

OpenStudy (anonymous):

I put this answer 1.4372 exactly how i got it

OpenStudy (mertsj):

Put the answer in like this: \[-\frac{9}{2}+\frac{\sqrt{141}}{2}\]

OpenStudy (mertsj):

When you give the decimal form of an irrational number you have to round it no matter how many decimal places you report because irrational numbers do not repeat of terminate and cannot be expressed exactly as a decimal.

OpenStudy (mertsj):

And you did round. I got 1.437171044 because my calculator displays 10 digits.

OpenStudy (anonymous):

even if i put that in its going to count it wrong because, its to long.

OpenStudy (mertsj):

I know. That's why I told you to enter the EXACT form

OpenStudy (anonymous):

so you want me to enter the fraction

OpenStudy (mertsj):

This is what I want you to enter:

OpenStudy (mertsj):

\[-\frac{9}{2}+\frac{\sqrt{141}}{2}\]

OpenStudy (mertsj):

I should have mentioned that before.

OpenStudy (anonymous):

uhm I can't enter that because of the sign my keyboard doesn't have it

OpenStudy (mertsj):

Typically those quizzes will have a way to enter a square root.

OpenStudy (mertsj):

Maybe the original problem is posted incorrectly. Go over it very carefully and make sure you haven't missed a sign or something.

OpenStudy (anonymous):

oh well I keep getting it wrong. and its right

OpenStudy (mertsj):

Can you copy and paste the original problem?

OpenStudy (anonymous):

The product of a number and the number increased by 9 is 15. What is the larger of the 2 solutions?

OpenStudy (anonymous):

it also says this when you get it wrong Translate to an equation. Get the equation equal to 0 before solving.

OpenStudy (mertsj):

Yes. It is giving you two hints to help you with the solution. But the question the problem asks is for the larger answer.

OpenStudy (mertsj):

Somewhere it shows you how to enter square roots.

OpenStudy (anonymous):

it doesnt show you anywhere!! I usually dont ever enter sqrts i enter decimals

OpenStudy (mertsj):

Did you try entering the answer to the nearest tenth?

OpenStudy (anonymous):

i already tried 1.44

OpenStudy (anonymous):

isnt that it to the nearest tenth

OpenStudy (mertsj):

Try 1.4

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

.......

OpenStudy (mertsj):

We know the answer is correct. Just have to find the correct way to enter it.

OpenStudy (anonymous):

okay. what else is there

OpenStudy (mertsj):

Did you ever find out?